OpenStudy (anonymous):

Add. -4/(x-1) + (1-3x)/x

OpenStudy (mathmale):

Hey, JayZee, What's the lowest common denominator of these two fractions? Hint: Multiply the first fraction (both numerator and denominator) by x. Multiply the second fraction (both numerator and denominator by ??? Multiply out terms in the numerator of the 2nd fraction. Then combine the numerators and write their sum over the lowest common denominator. Clear? Good. Not so clear? Please ask questions. Good luck!

OpenStudy (anonymous):

Ok, im about to try it out than check back with you on my answer

OpenStudy (studygurl14):

\[\frac{ -4 }{ x - 1 } + \frac{ 1 - 3x }{ x }\] \[\frac{ -4(x) }{ (x - 1)x } + \frac{ (1 - 3x)(x - 1) }{ x(x - 1) }\] \[\frac{ -4x + (1 - 3x)(x - 1) }{ x(x - 1) }\] \[\frac{ -4x + x - 1 - 3x^{2} + 3x}{ x(x - 1) }\] Can you do the rest?

OpenStudy (anonymous):

yeah thanks so much, do i leave the denominator the same or do i distrubute?

OpenStudy (mathmale):

I'd leave it as is. It's simpler that way, and there's no benefit from multiplying it out. I would, however, recommend that you combine like terms in the numerator. Good going.